# An active learning method for solving competitive multi-agent decision-making and control problems

### v3

@@ I-B Related work
Partially unrelated to the contributions mentioned so far, [27] addressed payoff-function learning as a standard regression problem. The methodology presented there, however, focused on games in normal form and came with no theoretical guarantees. Combining proximal-point iterations and ordinary least-squares (LS) estimators, [28] designed a distributed algorithm with probabilistic convergence guarantees to an equilibrium in stochastic games where the agents learn their own cost functions. In [29, 30], instead, a coordinator aims at reconstructing private information held by the agents to enable the distributed computation of an equilibrium by designing personalized incentives affecting the cost functions.
@@ I-C Summary of contributions and paper organization
The proofs of the technical results of the paper are all deferred to Appendix A, while Appendix A-B reports mathematical derivations adopted in the implementation of our approach.
@@
$\mathbb{N}$, $\mathbb{R}$ and $\mathbb{R}_{\geq 0}$ denote the set of natural, real, and nonnegative real numbers, respectively. $\mathbb{N}_{0}\coloneqq\mathbb{N}\cup\{0\}$, while $\bar{\mathbb{R}}\coloneqq\mathbb{R}\cup\{+\infty\}$ identifies the set of extended real numbers. Given a matrix $A\in\mathbb{R}^{n\times n}$, $A^{\top}$ denotes its transpose, while $A\succ 0$ ($\succcurlyeq 0$) denotes that $A$ is a positive (semi)definite matrix. Given a vector $v\in\mathbb{R}^{n}$ and a matrix $A\succ 0$, we denote by $\|v\|_{2}$ the standard Euclidean norm, while by $\|\cdot\|_{A}$ the $A$–induced norm $\|v\|_{A}\coloneqq\sqrt{v^{\top}Av}=\sqrt{\langle Av,v\rangle}$, where $\langle\cdot,\cdot\rangle:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ stands for the standard inner product. $\mathcal{B}_{\theta}(\bar{x})$ represents the $n$-dimensional ball centred around $\bar{x}$ with radius $\theta>0$, i.e., $\mathcal{B}_{\theta}(\bar{x})\coloneqq\left\{x\in\mathbb{R}^{n}\ \left|\ \|x-\bar{x}\|_{2}\leq\theta\right.\right\}$. $I_{n}$, $\boldsymbol{1}_{n}$, and $\boldsymbol{0}_{n}$ denote the $n\times n$ identity matrix, the vector of all $1$, and 00, respectively (we omit the dimension $n$ whenever clear). Given a set $\mathcal{X}\subseteq\mathbb{R}^{n}$, $\iota_{\mathcal{X}}:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ denote the associated indicator function, i.e., $\iota_{\mathcal{X}}(x)=0$ if $x\in\mathcal{X}$, $\iota_{\mathcal{X}}(x)=+\infty$ otherwise. The operator $\mathrm{col}(\cdot)$ stacks its arguments in column vectors or matrices of compatible dimensions. For example, given vectors $x_{1},\dots,x_{N}$ with $x_{i}\in\mathbb{R}^{n_{i}}$ and $\mathcal{N}=\{1,\dots,N\}$, we denote $\boldsymbol{x}\coloneqq(x_{1}^{\top},\dots,x_{N}^{\top})^{\top}=\mathrm{col}((x_{i})_{i\in\mathcal{N}})\in\mathbb{R}^{n}$, $n\coloneqq\sum_{i\in\mathcal{N}}n_{i}$, and $\boldsymbol{x}_{-i}\coloneqq\mathrm{col}((x_{j})_{j\in\mathcal{N}\setminus\{i\}})$. With a slight abuse of notation, we sometimes use also $\boldsymbol{x}=(x_{i},\boldsymbol{x}_{-i})$. The uniform distribution on the interval $[a,b]$ is denoted by $\mathcal{U}(a,b)$.
@@ II Learning problem
We now formalize the learning problem addressed, which is successively motivated through two multi-agent control and decision-making problems that fit the framework we investigate.
@@ III Preliminary results
Then, consider a generic function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\to\mathbb{R}$, which we evaluate as $f(x,\theta)$ with $x\in\mathbb{R}^{n}$ and $\theta\in\mathbb{R}^{p}$, and some feasible set $\mathcal{X}\subseteq\mathbb{R}^{n}$ satisfying the following conditions:
@@
The set $\mathcal{X}$ is convex. Moreover, for all $\theta\in\mathbb{R}^{p}$, $x\mapsto f(x,\theta)$ is a convex function. $\square$
@@
    For all $x\in\mathcal{X}$, $\theta\mapsto f(x,\theta)$ is convex and differentiable;
@@
    For all $\theta\in\mathbb{R}^{p}$, $x\mapsto f(x,\theta)$ is continuous;
@@
    For all $\theta\in\mathbb{R}^{p}$, the vector $\partial f(x,\theta)/\partial\theta\in\mathbb{R}^{p}$ of partial derivatives is bounded w.r.t. $x$. $\square$
@@
Let Assumption 3.1 hold true. Then, given any $\theta\in\mathbb{R}^{p}$, the set $\mathcal{M}(\theta)\coloneqq{\textnormal{{argmin}}}_{y\in\mathcal{X}}~{}f(y,\theta)$ is convex. If also Assumption 3.2 holds true, then the value function $f^{\star}:\mathbb{R}^{p}\to\mathbb{R}$, $f^{\star}(\theta)\coloneqq\textnormal{{min}}_{y\in\mathcal{X}}~{}f(y,\theta)$, is continuous. $\square$
@@
A function $h:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ is lower semicontinuous if, for all $\tilde{x}\in\mathbb{R}^{n}$,
$\lim_{x\rightarrow\tilde{x}}\inf h(x)\coloneqq\lim_{\varepsilon\rightarrow 0}\left(\inf_{y\in\mathcal{B}_{\varepsilon}(\tilde{x})}h(y)\right)\geq h(\tilde{x}).$square
@@
Let Assumption 3.1 and 3.2 hold true. Moreover, let $\mathcal{X}$ be a closed set. Then, the function $\bar{f}:\mathbb{R}^{n}\times\mathbb{R}^{p}\to\bar{\mathbb{R}}$, defined as $\bar{f}(x,\theta)\coloneqq f(x,\theta)+\iota_{\mathcal{X}}(x)$ is lower semicontinuous and level-bounded in $x$ locally uniformly in $\theta$. $\square$
We conclude this section with the following result establishing the convergence of the sequence of minimizers for $f$ with minimum norm, $\{x^{k}\}_{k\in\mathbb{N}}$. Specifically, given some $\theta^{k}\in\mathbb{R}^{p}$ the corresponding element $x^{k}$ is defined as follows:
@@ IV Active learning procedure and main results
We now introduce and discuss a scheme in which an external entity attempts to converge to a stationary action profile by actively learning local surrogates of the private action-reaction mappings through an iterative sample collection procedure.
@@ IV-A Algorithm description
The main steps of our active learning scheme are summarized in Algorithm 1, where the black-filled bullets refer to the tasks the external observer is required to perform, while the empty bullet to the one performed by the agents in $\mathcal{N}$.
@@
which contains all collective profiles that are the closest (according to the squared Euclidean norm, although different metrics could be used) to a fixed point of each $\hat{f}_{i}(\cdot,\theta_{i}^{k+1})$. This step is motivated by the notion of stationary action profile in Definition 2.1, which coincides with a fixed point of the stack of the agents’ mappings, i.e., $x_{i}^{\star}=f_{i}(\boldsymbol{x}_{-i}^{\star})$ for all $i\in\mathcal{N}$. Indeed, if $\hat{f}_{i}$ were exactly equal to $f_{i}$ and the minimum in (7) were zero, any $\boldsymbol{x}^{\star}\in\mathcal{M}(\theta^{k+1})$ would be a stationary action profile. Note that, by referring to (7), $\theta^{k+1}$ identifies the whole collection of parameters $\{\theta_{i}^{k+1}\}_{i\in\mathcal{N}}$ characterizing the surrogate mappings, which at every iteration represent the argument of the corresponding parameter-to-query mapping $\mathcal{M}(\cdot)$. In the spirit of the results developed in §III, the cost function in (7) then takes on the same role as the generic function $f$ in that section, thus establishing an immediate connection between (7) and the point-to-set mapping $\mathcal{M}(\cdot)$ defined in Lemma 3.3.
@@ IV-B Initialization
is solved for $k=1,\ldots,K_{\textrm{in}}$ to define $\boldsymbol{x}^{k}$ and pass it in parallel to each agent, which in turn responds through the private action-reaction mapping $f_{i}$. The resulting actions are collected by the external entity to update its estimates $\theta_{i}^{k}$. The remaining $K-K_{\textrm{in}}$ iterations of Algorithm 1 are then executed, where $K$ is the total number of queries performed to the agents.
Since such a random initialization procedure may affect the performance of the active learning scheme, we will discuss it in detail within each numerical example of §VI.
@@ IV-C Complexity analysis
In (12) $\beta\geq 0$ defines the learning rate of the filter, and the initial value of the $N$-by-$N$ covariance matrix $P_{i}^{0}=\alpha I$ the amount of $\ell_{2}$-regularization on $\theta_{i}$ (cf. [42, Eq. (12)]). In particular, the larger the $\beta$, the faster is the convergence of the linear estimator (we obtain standard recursive LS updates for $\beta=0$ [38, 39]), while the smaller the $\alpha$, the larger is the $\ell_{2}$-regularization. We will adopt the above setting in all the numerical simulations reported in §VI.
@@ IV-D Asymptotic properties
As a consequence of Theorem 4.5, in case the convergence of the parametric estimates happens, it can only be towards the true (at least pointwise) values. This fact, together with $\lim_{k\to\infty}\|\hat{\boldsymbol{x}}^{k}-\boldsymbol{x}^{k}\|_{2}=0$, yields $\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|^{2}_{2}=\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-f_{i}(\hat{\boldsymbol{x}}_{-i})\|^{2}_{2}=0$, and hence $\|\hat{\boldsymbol{x}}_{i}-\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|_{2}$ for all $i\in\mathcal{N}$, i.e., the external entity has in its hands both a stationary action profile $\hat{\boldsymbol{x}}$, and pointwise-exact surrogates of the action-reaction mappings $\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})=f_{i}(\hat{\boldsymbol{x}}_{-i})$.
@@ VI Numerical results
We now test the practical effectiveness of the active learning scheme in Algorithm 1 on several numerical instances of the multi-agent control and decision-making problems described in §I-A, §II, along with comparing its performance with the state-of-the-art centralized method for GNE computation in [43]. We adopt the affine setting (1) with MSE loss (11) detailed in §IV-C. All simulations are run in MATLAB on a laptop with an Apple M2 chip featuring an 8-core CPU and 16 Gb RAM.
@@ VI-B Generalized Nash games
Specifically, [45, Ex. 1] considers $N=10$ agents with cost functions $J_{i}(x_{i},\boldsymbol{x}_{-i})=N(1+(i-1)/2)x_{i}-x_{i}(60N-\boldsymbol{1}^{\top}_{N}\boldsymbol{x})$, while $\Omega=[7,100]^{N}$.
Figure 5: Hyperparameters analysis on the Nash equilibrium problem formalized in [45, Ex. 1].
@@
Figure 6: Nash equilibrium problem formalized in [45, Ex. 1]. The shaded regions correspond to the random data collection phase. Top: sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to an equilibrium (dashed lines). Bottom: convergent behaviour of $\{\theta_{i}^{k}\}_{k\in\mathbb{N}}$ to reconstruct the BR mapping of agent $1$. Only the first 25 samples out of the allocated $K=100$ are reported.
@@
Figure 7: GNEP formalized in [46, Ex. 1]. The shaded regions correspond to the random data collection phase. Top: sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to an equilibrium (dashed lines). Bottom: convergent behaviour of the parameter vector reconstructing the BR mapping of agent $1$.
Figure 8: GNEP described in [46, Ex. 3]. The shaded region corresponds to the data collection phase ($K_{\textrm{in}}=30$). Sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to a GNE (dashed lines).
Similar conclusions can be also drawn from Fig. 7 that illustrates the case in which Algorithm 1 is applied to the example described in [46, Ex. 1] (in this case, $K=50$). Modelling the internet switching behaviour induced by selfish users, in this example each cost function reads as:
@@
Then, while the first agent’s strategy is constrained so that $x_{1}\in[0.3,0.5]$, for all $j\neq 1$ we have $x_{j}\in[0.01,100]$. In addition, a standard linear constraint $\boldsymbol{1}^{\top}_{N}\boldsymbol{x}\leq 1$ couples the strategies of the whole population of decision-makers, thus actually resulting into a GNEP. An example with multidimensional decision vectors is instead the one illustrated in Fig. 8, which amounts to the GNEP described in [46, Ex. 3] and involving $N=3$ agents with three, two and two variables, respectively. Even though the point $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ converges to is different from the GNE reported in [46, Ex. 3], it still coincides with an equilibrium of the underlying GNEP, being a fixed point of the BR mappings.
Figure 9: GNEP formalized in [46, Ex. 11]. The shaded region corresponds to the random data collection phase ($K=100$). Top: Sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to the normalized equilibrium (dashed lines). Bottom: convergent behaviour of the parameter vectors reconstructing the BR mapping for both agents (same colour, solid lines correspond to the first element of each $\theta_{i}$, while dashed lines to the second).
@@
We conclude this part by discussing two simple, yet significant, examples involving two agents only. The first one consists in the GNEP in [46, Ex. 11] featuring an infinite number of equilibria $\left\{\mathrm{col}(\alpha,1-\alpha)\ \left|\ \alpha\in[1/2,1]\right.\right\}$, for which Fig. 9 shows that Algorithm 1 is actually able to return one of those solutions – specifically that coinciding with $\alpha=0.7$. After executing such an example $1000$ times, we have noticed that Algorithm 1 always returns the same equilibrium corresponding to $\alpha=0.7$ in all those instances in which it has converged. On the contrary, the second example considered does not have any Nash equilibrium due to the lack of quasiconvexity in the agents’ cost functions, which are taken as $J_{1}(x_{1},x_{2})=-x_{1}^{2}+2x_{1}x_{2}$ and $J_{2}=x_{2}-2x_{1}x_{2}$, $x_{i}\in[0,1]$ for $i\in\{1,2\}$. Figure 10 reports the numerical results for this example, where it is evident the non-convergent behaviour of the main quantities involved, particularly the parameter vectors $\theta_{i}$. According to the discussion in §V-A, in this case the sufficient conditions established in Theorem 4.5 do not hold, and therefore no certificate for the existence of Nash equilibria is provided, which is indeed the case. Finally, Table II shows that, in all those numerical instances considered for which convergence of the parametric estimates happens, also the sufficient condition for the uniqueness of the solution in $\mathcal{M}(\tilde{\theta})$ is met (namely, positive definiteness of the Hessian matrix $H$ of the cost function, along with convexity of $\Omega$), thus verifying the requirements in Theorem 4.5 numerically.

### v4

@@ I-B Related work
Suitable examples of works tackling standard problems in game theory through machine learning techniques can be found in [27, 28, 29, 30]. In particular, [27] addressed payoff-function learning as a standard regression problem. The methodology presented there, however, focused on games in normal form and came with no theoretical guarantees. Combining proximal-point iterations and ordinary least-squares (LS) estimators, [28] designed a distributed algorithm with probabilistic convergence guarantees to an equilibrium in stochastic games where the agents learn their own cost functions. In [29, 30], instead, a coordinator aims at reconstructing private information held by the agents to enable the computation of an equilibrium by designing personalized incentives affecting the cost functions.
@@ I-C Summary of contributions and paper organization
The proofs of the technical results of the paper are all deferred to Appendix A, while Appendix A-B reports mathematical derivations adopted in the implementation of our approach. The Python package gnep-learn, implementing the proposed active learning-based scheme, is publicly available at https://github.com/bemporad/gnep-learn.
@@
$\mathbb{N}$, $\mathbb{R}$ and $\mathbb{R}_{\geq 0}$ denote the set of natural, real, and nonnegative real numbers, respectively. $\mathbb{N}_{0}\coloneqq\mathbb{N}\cup\{0\}$, while $\bar{\mathbb{R}}\coloneqq\mathbb{R}\cup\{+\infty\}$. The transpose of a matrix $A\in\mathbb{R}^{n\times n}$ is $A^{\top}$, while $A\succ 0$ ($\succcurlyeq 0$) denotes its positive (semi)definiteness. For a vector $v\in\mathbb{R}^{n}$ and a matrix $A\succ 0$, $\|v\|_{2}$ denotes the standard Euclidean norm, while $\|\cdot\|_{A}$ the $A$–induced norm $\|v\|_{A}\coloneqq\sqrt{v^{\top}Av}$. The $n$-dimensional ball centred around $\bar{x}$ with radius $\theta>0$ is $\mathcal{B}_{\theta}(\bar{x})\coloneqq\left\{x\in\mathbb{R}^{n}\ \left|\ \|x-\bar{x}\|_{2}\leq\theta\right.\right\}$. $I_{n}$, $\boldsymbol{1}_{n}$, and $\boldsymbol{0}_{n}$ denote the $n\times n$ identity matrix, the vector of all $1$, and 00, respectively (we omit the dimension $n$ whenever clear). Given a set $\mathcal{X}\subseteq\mathbb{R}^{n}$, $\iota_{\mathcal{X}}:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ denote the associated indicator function, i.e., $\iota_{\mathcal{X}}(x)=0$ if $x\in\mathcal{X}$, $\iota_{\mathcal{X}}(x)=+\infty$ otherwise. The operator $\mathrm{col}(\cdot)$ stacks its arguments in column vectors or matrices of compatible dimensions. For example, given vectors $x_{1},\dots,x_{N}$ with $x_{i}\in\mathbb{R}^{n_{i}}$ and $\mathcal{N}=\{1,\dots,N\}$, we denote $\boldsymbol{x}\coloneqq(x_{1}^{\top},\dots,x_{N}^{\top})^{\top}=\mathrm{col}((x_{i})_{i\in\mathcal{N}})\in\mathbb{R}^{n}$, $n\coloneqq\sum_{i\in\mathcal{N}}n_{i}$, and $\boldsymbol{x}_{-i}\coloneqq\mathrm{col}((x_{j})_{j\in\mathcal{N}\setminus\{i\}})$. With a slight abuse of notation, we sometimes use also $\boldsymbol{x}=(x_{i},\boldsymbol{x}_{-i})$. The uniform distribution on the interval $[a,b]$ is denoted by $\mathcal{U}(a,b)$.
@@ II Learning problem
We start by formalizing the learning problem addressed, which is successively motivated by two multi-agent control and decision-making problems that fit the framework we consider.
@@ III Preliminary results
Let us then consider a generic function $r:\mathbb{R}^{n}\times\mathbb{R}^{p}\to\mathbb{R}$, evaluated as $r(x,\theta)$ with $x\in\mathbb{R}^{n}$ and $\theta\in\mathbb{R}^{p}$, and a feasible set $\mathcal{X}\subseteq\mathbb{R}^{n}$ satisfying the following conditions:
@@
The set $\mathcal{X}$ is convex. Moreover, for all $\theta\in\mathbb{R}^{p}$, $x\mapsto r(x,\theta)$ is a convex function. $\square$
@@
    For all $x\in\mathcal{X}$, $\theta\mapsto r(x,\theta)$ is convex and differentiable;
@@
    For all $\theta\in\mathbb{R}^{p}$, $x\mapsto r(x,\theta)$ is continuous;
@@
    For all $\theta\in\mathbb{R}^{p}$, the vector $\partial r(x,\theta)/\partial\theta\in\mathbb{R}^{p}$ of partial derivatives is bounded w.r.t. $x$. $\square$
@@
Let Assumption 3.1 hold true. Then, given any $\theta\in\mathbb{R}^{p}$, the set $\mathcal{M}(\theta)\coloneqq{\textnormal{{argmin}}}_{y\in\mathcal{X}}~{}r(y,\theta)$ is convex. If also Assumption 3.2 holds true, then the value function $r^{\star}:\mathbb{R}^{p}\to\mathbb{R}$, $r^{\star}(\theta)\coloneqq\textnormal{{min}}_{y\in\mathcal{X}}~{}r(y,\theta)$, is continuous. $\square$
@@
A function $h:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ is lower semicontinuous if, for all $\tilde{x}\in\mathbb{R}^{n}$, $\underset{x\rightarrow\tilde{x}}{\lim}\inf h(x)\coloneqq\underset{\varepsilon\rightarrow 0}{\lim}\left(\underset{y\in\mathcal{B}_{\varepsilon}(\tilde{x})}{\inf}h(y)\right)\geq h(\tilde{x})$. $\square$
@@
Let Assumption 3.1 and 3.2 hold true. Moreover, let $\mathcal{X}$ be a closed set. Then, the function $\bar{r}:\mathbb{R}^{n}\times\mathbb{R}^{p}\to\bar{\mathbb{R}}$, defined as $\bar{r}(x,\theta)\coloneqq r(x,\theta)+\iota_{\mathcal{X}}(x)$ is lower semicontinuous and level-bounded in $x$ locally uniformly in $\theta$. $\square$
We conclude this section with the following result establishing the convergence of the sequence of minimizers for $r$ with minimum norm, $\{x^{k}\}_{k\in\mathbb{N}}$. Specifically, given some $\theta^{k}\in\mathbb{R}^{p}$ the corresponding element $x^{k}$ is defined as follows:
@@ IV Active learning procedure and main results
We now introduce, discuss and analyze the convergence property of our iterative scheme based on a active learning.
@@ IV-A Algorithm description
The main steps of our active learning scheme are summarized in Algorithm $1$A Python package with its implementation is publicly available at https://github.com/bemporad/gnep-learn., where the black-filled bullets refer to the tasks the external observer is required to perform, while the empty bullet to the one performed by the agents in $\mathcal{N}$.
@@
which contains all collective profiles that are the closest (according to the squared Euclidean norm, although different metrics could be used) to a fixed point of each $\hat{f}_{i}(\cdot,\theta_{i}^{k+1})$. This step is motivated by the notion of stationary action profile in Definition 2.1, which coincides with a fixed point of the stack of the agents’ mappings, i.e., $x_{i}^{\star}=f_{i}(\boldsymbol{x}_{-i}^{\star})$ for all $i\in\mathcal{N}$. Indeed, if $\hat{f}_{i}$ were exactly equal to $f_{i}$ and the minimum in (7) were zero, any $\boldsymbol{x}^{\star}\in\mathcal{M}(\theta^{k+1})$ would be a stationary action profile. Note that, by referring to (7), $\theta^{k+1}$ identifies the whole collection of parameters $\{\theta_{i}^{k+1}\}_{i\in\mathcal{N}}$ characterizing the surrogate mappings, which at every iteration represent the argument of the corresponding parameter-to-query mapping $\mathcal{M}(\cdot)$. In the spirit of the results developed in §III, the cost function in (7) then takes on the same role as the generic function $r(\cdot)$ in that section, thus establishing an immediate connection between (7) and the point-to-set mapping $\mathcal{M}(\cdot)$ defined in Lemma 3.3.
@@ IV-B Initialization
is solved for $k=1,\ldots,K_{\textrm{in}}$ to define $\boldsymbol{x}^{k}$ and pass it in parallel to each agent, which in turn responds through the private action-reaction mapping $f_{i}$. The resulting actions are collected by the external entity to update its estimates $\theta_{i}^{k}$. The remaining $K-K_{\textrm{in}}$ iterations of Algorithm 1 are then executed, where $K$ is the total number of queries performed to the agents. Alternatively, more sophisticated initialization strategies are possible, such as Latin hypercube sampling [38, 39] as in, e.g., [4], however, significant differences are not expected in our framework as we only learn linear proxies rather than global surrogates of the agents’ responses. Nevertheless, given its key role we will discuss in detail the random initialization procedure within each numerical example reported in §VI.
@@ IV-C Complexity analysis
In (12) $\beta\geq 0$ defines the learning rate of the filter, and the initial value of the $N$-by-$N$ covariance matrix $P_{i}^{0}=\alpha I$ the amount of $\ell_{2}$-regularization on $\theta_{i}$ (cf. [44, Eq. (12)]). In particular, the larger the $\beta$, the faster is the convergence of the linear estimator (we obtain standard recursive LS updates for $\beta=0$ [40, 41]), while the smaller the $\alpha$, the larger is the $\ell_{2}$-regularization. We will adopt the above setting in all the numerical simulations reported in §VI.
@@ IV-D Asymptotic properties
As a consequence of Theorem 4.5, in case the convergence of the parametric estimates happens, it can only be towards the true (at least pointwise) values. This fact, together with $\lim_{k\to\infty}\|\hat{\boldsymbol{x}}^{k}-\boldsymbol{x}^{k}\|_{2}=0$, yields $\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|^{2}_{2}=\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-f_{i}(\hat{\boldsymbol{x}}_{-i})\|^{2}_{2}=0$, and hence $\|\hat{\boldsymbol{x}}_{i}-\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|_{2}=0$ for all $i\in\mathcal{N}$, i.e., the external entity has in its hands both a stationary action profile $\hat{\boldsymbol{x}}$, and pointwise-exact surrogates of the action-reaction mappings $\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})=f_{i}(\hat{\boldsymbol{x}}_{-i})$.
@@ VI Numerical results
We now test the practical effectiveness of the active learning scheme in Algorithm 1 on several numerical instances of the multi-agent control and decision-making problems described in §I-A, §II, along with comparing its performance with the state-of-the-art centralized method for GNE computation in [45]. We adopt the affine setting (1) with MSE loss (11) detailed in §IV-C. All simulations are run in MATLAB on a laptop with an Apple M2 chip featuring an 8-core CPU and 16 GB RAM. Some of the examples analyzed in this section are also included in the Python package gnep-learn, publicly available at https://github.com/bemporad/gnep-learn.
@@ VI-B Generalized Nash games
Specifically, [47, Ex. 1] considers $N=10$ agents with cost functions $J_{i}(x_{i},\boldsymbol{x}_{-i})=N(1+(i-1)/2)x_{i}-x_{i}(60N-\boldsymbol{1}^{\top}_{N}\boldsymbol{x})$, while $\Omega=[7,100]^{N}$.
Figure 5: Hyperparameters analysis on the Nash equilibrium problem formalized in [47, Ex. 1].
@@
Figure 6: Nash equilibrium problem formalized in [47, Ex. 1]. The shaded regions correspond to the random data collection phase. Top: sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to an equilibrium (dashed lines). Bottom: convergent behaviour of $\{\theta_{i}^{k}\}_{k\in\mathbb{N}}$ to reconstruct the BR mapping of agent $1$. Only the first 25 samples out of the allocated $K=100$ are reported.
@@
Figure 7: GNEP formalized in [48, Ex. 1]. The shaded regions correspond to the random data collection phase. Top: sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to an equilibrium (dashed lines). Bottom: convergent behaviour of the parameter vector reconstructing the BR mapping of agent $1$.
Figure 8: GNEP described in [48, Ex. 3]. The shaded region corresponds to the data collection phase ($K_{\textrm{in}}=30$). Sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to a GNE (dashed lines).
Similar conclusions can be also drawn from Fig. 7 that illustrates the case in which Algorithm 1 is applied to the example described in [48, Ex. 1] (in this case, $K=50$). Modelling the internet switching behaviour induced by selfish users, in this example each cost function reads as:
@@
Then, while the first agent’s strategy is constrained so that $x_{1}\in[0.3,0.5]$, for all $j\neq 1$ we have $x_{j}\in[0.01,100]$. In addition, a standard linear constraint $\boldsymbol{1}^{\top}_{N}\boldsymbol{x}\leq 1$ couples the strategies of the whole population of decision-makers, thus actually resulting into a GNEP. An example with multidimensional decision vectors is instead the one illustrated in Fig. 8, which amounts to the GNEP described in [48, Ex. 3] and involving $N=3$ agents with three, two and two variables, respectively. Even though the point $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ converges to is different from the GNE reported in [48, Ex. 3], it still coincides with an equilibrium of the underlying GNEP, being a fixed point of the BR mappings.
Figure 9: GNEP formalized in [48, Ex. 11]. The shaded region corresponds to the random data collection phase ($K=100$). Top: Sequence $\{\hat{\boldsymbol{x}}^{k}\}_{k\in\mathbb{N}}$ (solid lines) converging to the normalized equilibrium (dashed lines). Bottom: convergent behaviour of the parameter vectors reconstructing the BR mapping for both agents (same colour, solid lines correspond to the first element of each $\theta_{i}$, while dashed lines to the second).
@@
We conclude this part by discussing two simple, yet significant, examples involving two agents only. The first one consists in the GNEP in [48, Ex. 11] featuring an infinite number of equilibria $\left\{\mathrm{col}(\alpha,1-\alpha)\ \left|\ \alpha\in[1/2,1]\right.\right\}$, for which Fig. 9 shows that Algorithm 1 is actually able to return one of those solutions – specifically that coinciding with $\alpha=0.7$. After executing such an example $1000$ times, we have noticed that Algorithm 1 always returns the same equilibrium corresponding to $\alpha=0.7$ in all those instances in which it has converged. On the contrary, the second example considered does not have any Nash equilibrium due to the lack of quasiconvexity in the agents’ cost functions, which are taken as $J_{1}(x_{1},x_{2})=-x_{1}^{2}+2x_{1}x_{2}$ and $J_{2}=x_{2}-2x_{1}x_{2}$, $x_{i}\in[0,1]$ for $i\in\{1,2\}$. Figure 10 reports the numerical results for this example, where it is evident the non-convergent behaviour of the main quantities involved, particularly the parameter vectors $\theta_{i}$. According to the discussion in §V-A, in this case the sufficient conditions established in Theorem 4.5 do not hold, and therefore no certificate for the existence of Nash equilibria is provided, which is indeed the case. Finally, Table II shows that, in all those numerical instances considered for which convergence of the parametric estimates happens, also the sufficient condition for the uniqueness of the solution in $\mathcal{M}(\tilde{\theta})$ is met (namely, positive definiteness of the Hessian matrix $H$ of the cost function, along with convexity of $\Omega$), thus verifying the requirements in Theorem 4.5 numerically.